A polystyrene, having formula Br3C6H3(C8H8)n was prepared by heating styrene with tribromobenzoyl peroxide in the abscense of air. If it was found to contain 10.46% bromine by weight. Find the value of n.


plz give complete solution.....

Grade:11

2 Answers

Sudheesh Singanamalla
13 years ago

Let the weight of polystyrene prepared be 100 g

number of moles of Br in 100g of ploystyrene = 10.46 / 79.9 = 0.1309

From the formula of polystyrene , we have ,

number of moles of Br = 3 * number of moles of Br3C6H3(C8H8)n

0.1309 = 3 * Wt / molar weight

0.1309 = 3*100 / (314.7 + 44n)

0.1309 * ( 314.7 + 44n ) = 300

(314.7 + 44n) = 2291.825

44n = 2291.825 - 314.7

n = 1977.125/44

n = 44.93

n = 45 (approx)

Therefore value of n is 45

-----------------------------------------------------------------------------------------------

Please Approve !

-----------------------------------------------------------

Rudransh Kumar Ankodia
5 years ago
100g polystyrene contains 10.46g Br
240g Br(1mole polystyrene ) 
By unitary method wecan find Molar mass of polystyrene 
(100*240)/10.46=2294.45
MM=2294.45
MM-Mass of each component = 0
2294.45-(240+72+3+104n)=0
So n=19.03
If you agree with this solution 
Plz approve